x+y+z=1800
x=2(y+z)
20x+35y+50z=4800

Answer by TimothyLamb(4379)

x+y+z=1800
x=2(y+z)
20x+35y+50z=4800
---
put the system of linear equations into standard form
---
x+y+z=1800
-x+2y+2z=0
20x+35y+50z=4800
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 1200
y = 3280
z = -2680
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---